OpenStudy (anonymous):

55/100 in lowest terms and decimal

OpenStudy (anonymous):

55/100 = 11/20 = 0.55

OpenStudy (anonymous):

0.55

OpenStudy (anonymous):

.55

OpenStudy (anonymous):

(55/5)/(100/5) = 11/20 = 0.55